Magnitude of vectors

Arrange the following vectors in the order of their magnitude, with the vector of largest magnitude first. (i) A = ( 3i + 5j - 2k) m; (ii) B = (-3i + 5j -2k) m; (iii)C = (3i - 5j - 2k ) m; (iv) D = (3i _ 5j +2k ) m.
Answer #1

The magnitude of a vector is given by the square root of the sum of the squares of its components. (basically the Pythagorean theorem in 3-D)
M=sqrt(x2+y2+z2)
so
i. M=sqrt(9+25+4)=sqrt(38)=6.16
ii. M=sqrt(9+25+4)=6.16
iii. M=sqrt(9+25+4)=6.16
iv. M=6.16 again
they are all equal
i=ii=iii=iv

answered by: Gottara